OpenStudy (anonymous):

Use equations to solve the following questions Daniel and Travis are twins, but their sister Phillipa is 3 years older. If the sum of their three ages is 36, how old are the twins?

OpenStudy (anonymous):

Let the age of Phillipa is : y: And the twins age be : x So: y = x + 3 So: x + x + y = 36 So: 2x + x + 3 = 36 3x = 33 Find x from here..

OpenStudy (anonymous):

Divide both the sides by 3and you will get x..

OpenStudy (phoenixfire):

Let the twins age be a and Phillipa be b total age is equal to sum of all ages or 36=a+a+b or 36=2a+b you know Phillipa is 3 years older, so b=a+3 substitute that back in: 36=2a+(a+3) 36=3a+3 Now solve for a to find the twins age. 36-3=3a 33/3=a 11=a Twins are 11 years old.

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

Welcome dear..